# Rethinking Diffusion Posterior Sampling: From Conditional Score Estimator to Maximizing a Posterior

### v1

@@ Abstract
Recent advancements in diffusion models have been leveraged to address inverse problems without additional training, and Diffusion Posterior Sampling (DPS) (Chung et al., 2022a) is among the most popular approaches. Previous analyses suggest that DPS accomplishes posterior sampling by approximating the conditional score. While in this paper, we demonstrate that the conditional score approximation employed by DPS is not as effective as previously assumed, but rather aligns more closely with the principle of maximizing a posterior (MAP). This assertion is substantiated through an examination of DPS on 512$\times$512 ImageNet images, revealing that: 1) DPS’s conditional score estimation significantly diverges from the score of a well-trained conditional diffusion model and is even inferior to the unconditional score; 2) The mean of DPS’s conditional score estimation deviates significantly from zero, rendering it an invalid score estimation; 3) DPS generates high-quality samples with significantly lower diversity. In light of the above findings, we posit that DPS more closely resembles MAP than a conditional score estimator, and accordingly propose the following enhancements to DPS: 1) we explicitly maximize the posterior through multi-step gradient ascent and projection; 2) we utilize a light-weighted conditional score estimator trained with only 100 images and 8 GPU hours. Extensive experimental results indicate that these proposed improvements significantly enhance DPS’s performance. The source code for these improvements is provided in https://github.com/tongdaxu/Rethinking-Diffusion-Posterior-Sampling-From-Conditional-Score-Estimator-to-Maximizing-a-Posterior.
@@ 1 Introduction
From our empirical study, we make three major observations: 1). The conditional score estimation of DPS considerably diverges from the score of a properly trained conditional diffusion model and is even outperformed by an unconditional score; 2). The mean of DPS’s conditional score estimate significantly deviates from zero, thus failing to qualify as a valid score estimation; 3). The samples generated by DPS, although of high quality, exhibit markedly lower diversity. These findings collectively argue that DPS more closely aligns with the MAP framework than with conditional score estimation.
@@ 2 Preliminaries
where $\beta_{t}$ are hyperparameters. The reverse process of DDPM is also a Markov chain, with the transition kernel $p_{\theta}(X_{t-1}|X_{t})$ depending on the score function $\nabla_{X_{t}}\log p(X_{t})$, which is approximated by a neural network $s_{\theta}(X_{t},t)$ parameterized by $\theta$:
@@ 3.2 Observation I: DPS Has Large Score Error
To verify that StableSR is a reliable baseline for conditional score, we train a super-resolution ControlNet (Zhang et al., 2023), which uses a completely different neural network from StableSR. However, as shown in Figure 1 the score error of the ControlNet is significantly smaller than that of the unconditional score. And in Table 1, we show that StableSR outperforms DPS in all perceptual metrics, such as FID (Fréchet inception distance) (Heusel et al., 2017), KID (Kernel Inception Distance) (Binkowski et al., 2018) and LPIPS (Learned Perceptual Image Patch Similarity) (Zhang et al., 2018), confirming that it is a reliable estimate of the conditional score.
@@ 3.4 Observation III: DPS Has Low Sample Diversity
Following the previous SR×$\times 8$ settings, we draw $k=50$ samples from DPS and StableSR respectively and compute the per-pixel standard deviation. The results are shown in Figure 3 and Table 3. We observe that DPS with $\zeta_{t}=4.8$ has a much lower per-pixel standard deviation compared to StableSR. In Figure 4, we show that the samples from StableSR contain more visually meaningful variations than those from DPS. This phenomenon is also observed by Cohen et al. (2023).
@@ 3.5 Hypothesis: DPS Aligns More Closely to Maximizing a Posterior
With Observation I & II, one can conclude that DPS is not an effective conditional score estimator for image restoration. The score error of DPS being even larger than the unconditional score, along with its high score mean indicating it is not a proper score, supports this conclusion.
However, DPS does work well and produces high-quality samples, despite being a poor conditional score estimator. Along with Observation III that DPS has low sample diversity, we hypothesize that DPS is closer to another paradigm of image restoration other than posterior sampling: the maximization of a posterior (MAP) estimate.
@@
Besides, the MAP hypothesis also explains previous works that are not explainable with conditional score estimation theory, such as why DPS works with a high score error lower-bound (Yang et al., 2024), why Adam (Kingma & Ba, 2014) helps DPS (He et al., 2023; Chung et al., 2023), and why DPS has low sample diversity (Cohen et al., 2023) (See Appendix C.1).
@@ 4.2 Improvement II: A Light-weight Conditional Score Estimator
In practice, we adopt ControlNet (Zhang et al., 2023) as our CSE, which is a widely used neural network architecture for adding control to Stable Diffusion. We are surprised to find that our CSE is very efficient in terms of both data and temporal complexity. Specifically, it only takes 8 GPU hours and 100 images to train, which is roughly equivalent to the time of running DPS for 100 images. For DPS evaluated with 1000 images, the additional training time of CSE is around $28$s for each image, while the DPS itself costs several minutes. Alternatively, we can also use fully synthetic images generated by Stable Diffusion, in which case our CSE does not require any new data.
@@ 5.1 Experimental Setup
Dataset & Diffusion Model We utilize the first 1000 images from the ImageNet validation split. All images are resized such that their shorter edge is 512 pixels, and then center-cropped to dimensions of $512\times 512$ pixels. We use Stable Diffusion 2.0 as the base diffusion model and employ a 500-step ancestral sampling solver (i.e., DDPM) to align with our assumptions. There are some works using stochastic ancestral sampling solvers (Yu et al., 2023; Yang et al., 2024) while some other works (Chung et al., 2023; Rout et al., 2023) using deterministic PF-ODE solvers (i.e., DDIM). We discuss the impact of base diffusion models, solvers and detailed setup in Appendix. B.4.
@@ 7 Discussion & Conclusion
One limitation of the conditional score estimator used in Improvement II is that it employs an established approach, ControlNet (Zhang et al., 2023), which is not optimized for sample and computational efficiency. Currently, we can improve DPS with 100 images and 8 GPU hours using the vanilla ControlNet. However, we believe that more work on neural network architecture and training strategies could further reduce the number of required images and the training time. It would be particularly interesting if we could improve DPS using just a handful of examples and shorter time, extending its applicability from zero-shot to few-shot scenarios.
To conclude, we demonstrate that DPS does not function as a conditional score estimator in real-life scenarios through three key observations. Instead, we hypothesize that DPS is closer to maximizing a posterior distribution. Based on this new hypothesis, we propose two improvements to DPS. Empirical results show that both of our proposed improvements significantly enhance the performance of DPS.

### v2

@@ Abstract
Recent advancements in diffusion models have been leveraged to address inverse problems without additional training, and Diffusion Posterior Sampling (DPS) (Chung et al., 2022a) is among the most popular approaches. Previous analyses suggest that DPS accomplishes posterior sampling by approximating the conditional score. While in this paper, we demonstrate that the conditional score approximation employed by DPS is not as effective as previously assumed, but rather aligns more closely with the principle of maximizing a posterior (MAP). This assertion is substantiated through an examination of DPS on 512$\times$512 ImageNet images, revealing that: 1) DPS’s conditional score estimation significantly diverges from the score of a well-trained conditional diffusion model and is even inferior to the unconditional score; 2) The mean of DPS’s conditional score estimation deviates significantly from zero, rendering it an invalid score estimation; 3) DPS generates high-quality samples with significantly lower diversity. In light of the above findings, we posit that DPS more closely resembles MAP than a conditional score estimator, and accordingly propose the following enhancements to DPS: 1) we explicitly maximize the posterior through multi-step gradient ascent and projection; 2) we utilize a light-weighted conditional score estimator trained with only 100 images and 8 GPU hours. Extensive experimental results indicate that these proposed improvements significantly enhance DPS’s performance. The source code for these improvements is provided in this link.
@@ 1 Introduction
From our empirical study, we make three major observations: 1). The conditional score estimation of DPS considerably diverges from the score of a properly trained conditional diffusion model and is even outperformed by an unconditional score; 2). The mean of DPS’s conditional score estimate significantly deviates from zero, thus failing to qualify as a valid score; 3). The samples generated by DPS, although of high quality, exhibit markedly lower diversity. These findings collectively argue that DPS more closely aligns with the MAP framework than with conditional score estimation.
@@ 2 Preliminaries
where $\beta_{t}$ are hyperparameters. The reverse process of DDPM is a Markov chain with the transition $p_{\theta}(X_{t-1}|X_{t})$, with the score function $\nabla_{X_{t}}\log p(X_{t})$ approximated by a neural network $s_{\theta}(X_{t},t)$:
@@ 3.2 Observation I: DPS Has Large Score Error
To verify that StableSR is a reliable baseline for conditional score, we train a super-resolution ControlNet (Zhang et al., 2023), which uses a different neural network. However, as shown in Figure 1 the score error of the ControlNet is significantly smaller than unconditional score. Besides, StableSR outperforms DPS in FID (Fréchet inception distance) (Heusel et al., 2017), KID (Kernel Inception Distance) (Binkowski et al., 2018) and LPIPS (Learned Perceptual Image Patch Similarity) (Zhang et al., 2018), confirming that it is a reliable estimate of the conditional score.
@@ 3.4 Observation III: DPS Has Low Sample Diversity
We draw $k=50$ samples from DPS and StableSR respectively and compute the per-pixel standard deviation. The results are shown in Figure 3 and Table 3. We observe that DPS with $\zeta_{t}=4.8$ has a much lower standard deviation than StableSR. In Figure 4, we show that the samples from StableSR contain more visual variations than DPS, which is also observed by Cohen et al. (2023).
@@ 3.5 Hypothesis: DPS Aligns More Closely to Maximizing a Posterior
With Observation I & II, one can conclude that DPS is not an effective conditional score estimator for image restoration. However, DPS does work well and produces high-quality samples, despite being a poor conditional score estimator. Along with Observation III that DPS has low sample diversity, we hypothesize that DPS is closer to another paradigm of image restoration other than posterior sampling: the maximization of a posterior (MAP) estimate.
@@
Besides, the MAP hypothesis also explains previous works that are not explainable with conditional score estimation theory, such as why DPS has a high score error lower-bound (Yang et al., 2024) and why Adam helps DPS (He et al., 2023; Chung et al., 2023) (See Appendix C.1).
@@ 4.2 Improvement II: A Light-weight Conditional Score Estimator
In practice, we adopt ControlNet (Zhang et al., 2023) as CSE. We are surprised to find that our CSE is very efficient in terms of both data and temporal complexity. Specifically, it only takes 8 GPU hours and 100 images to train, which is roughly equivalent to the time of running DPS for 100 images. For DPS evaluated with 1000 images, the additional training time of CSE is around $28$s for each image, while the DPS itself costs several minutes. Alternatively, we can also use fully synthetic images generated by Stable Diffusion, in which case our CSE does not require any new data.
@@ 5.1 Experimental Setup
Dataset & Diffusion Model We utilize the first 1000 images from the ImageNet validation split. All images are resized and center-cropped to $512^{2}$ pixels. We use Stable Diffusion 2.0 as the base diffusion model and employ a 500-step ancestral sampling solver (i.e., DDPM) to align with our assumptions. There are some works using ancestral sampling solvers (Yu et al., 2023; Yang et al., 2024) while some other works (Chung et al., 2023; Rout et al., 2023) using PF-ODE solvers (i.e., DDIM). We discuss the impact of base diffusion models, solvers and detailed setup in Appendix. B.4.
@@ 6 Related Work
During the conference, the authors becomes aware of Gutha et al. (2024), a recent work that also views DPS as a MAP problem and use multi-step gradient ascent. While this work does not consider the theoretical formulation and does not employ projection between gradient ascent steps.
@@ 7 Discussion & Conclusion
One limitation of the conditional score estimator used in Improvement II is that it employs ControlNet (Zhang et al., 2023), which is not optimized for sample and computational efficiency. It would be particularly interesting if we could improve DPS using just a handful of examples and shorter time, extending its applicability from zero-shot to few-shot scenarios.
To conclude, we demonstrate that DPS does not function as a conditional score estimator through three key observations. Instead, we hypothesize that DPS is closer to maximizing a posterior distribution. Based on this new hypothesis, we propose two improvements to DPS. Empirical results show that both of our proposed improvements significantly enhance the performance of DPS.
Acknowledgments
This work is supported by Wuxi Research Institute of Applied Technologies, Tsinghua University under Grant 20242001120. No Patent or IP Claims.